Solve the equation:
1/20v - 1/40 + 1/4v = 1/4 + v

Answer by checkley71(8403)

1/20V-1/40+1/4V=1/4+V
1/20V+1/4V-V=1/4+1/40
1/20V+5/20V-20/20V=10/40+1/40
-14/20V=11/40
-14V*40=20*11
-560V=220
V=220/-560
V=-22/56
V=-11/28